# TabTreeFormer: Tabular Data Generation Using Hybrid Tree-Transformer

### v2

@@ 4.5 Efficiency
Compared to baseline auto-regressive transformers, TabTreeFormer demonstrates significant advantage in terms of efficiency, especially in training time on large datasets (see Table 5) and in generation time on datasets with a large number of features (see Fig. 9), which is consistent with our expectation due to the smaller model size and sampling without rejection. Raw training and generation time records can be found in Table 13 in Appendix D.5.
@@ 4.6 Ablation Study
Table 6: Ablation studies. The values show the relative score change when certain component in TabTreeFormer-S is taken out.
@@
On each core components of TabTreeFormer, including tree-based model (TBM), dual-quantization tokenization (DQT), and OCEL, we apply ablation study using TabTreeFormer-S on 2 randomly selected datasets: credit and diabetes, evaluated using MLE. The results, presented in Table 6 with raw values presented in Table 14 in Appendix D.6, verify the usefulness of all the core components.
@@ 5 Conclusion
In this paper, we propose TabTreeFormer, a hybrid tree-transformer for generating high-quality tabular data. TabTreeFormer inherits unique inductive biases from a tree-based model and also captures multimodal distribution to optimize the data representations of numeric values via a dual-quantization tokenizer. We comprehensively evaluate the performance of TabTreeFormer and show that it demonstrates superior fidelity, utility, privacy and outstanding efficiency with significantly smaller model size and computation time.
@@ A.2 Proof of Theorem 1
Suppose we have $\mathbf{z}$ and $\mathbf{z}^{*}$ as logits on vocabulary size $V$ and target label $t$, which satisfy both the following constraints:
  • For some $i\neq j\in\{1,\dots,V\}\setminus$ and $|t-i|\neq|t-j|$, $a=z_{i}=z_{j}^{*}\neq z_{j}=z_{i}^{*}=b$ for some $a,b\in\mathbb{R}$.
  • $z_{k}=z_{k}^{*},\forall k\in\{0,1,\dots,V-1\}\setminus\{i,j,t\}$.
Without loss of generality, suppose $d_{1}=|t-i|<|t-j|=d_{2},a<b$. Note that $w_{k}$ is monotonically increasing with $|t-k|$.
Thus, $e^{a}<e^{b},w_{ti}<w_{tj}$. This gives $\left(w_{ti}-w_{tj}\right)\left(e^{a}-e^{b}\right)<0$, so $P=w_{ti}e^{z_{i}}+w_{tj}e^{z_{j}}=w_{ti}e^{a}+w_{tj}e^{b}<w_{ti}e^{z_{i}^{*}}+w_{tj}e^{z_{j}^{*}}=w_{ti}e^{b}+w_{tj}e^{a}=Q$ for some $P,Q\in\mathbb{R}$.
Substitute into $\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ and $\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$,
$\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ $=-\log\frac{w_{tt}e^{z_{t}}}{P+\sum_{k\neq i,j}w_{k}e^{z_{k}}}$
$\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$ $=-\log\frac{w_{tt}e^{z_{t}}}{Q+\sum_{k\neq i,j}w_{k}e^{z_{k}}}$
Hence, $\mathcal{L}_{\text{oce}}(\mathbf{z},t)<\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$. The logits with higher logit value at position closer to target token is penalized less. ∎
@@ C.3 Baseline Implementation
We use the Synthcity (Qian et al., 2024) implementation for ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), GReaT (Borisov et al., 2023), TabDDPM (Kotelnikov et al., 2023), and TVAE (Xu et al., 2019). We also include CTAB-GAN+ (Zhao et al., 2024), REaLTabFormer (Solatorio & Dupriez, 2023), and TabSyn (Zhang et al., 2024) using their official code on GitHub or public SDK.
@@ D.1 Supplementary Fidelity Results
Table 9: Performance comparison between different models in terms of fidelity. The closer the values are to 1, the better the fidelity is. Best scores and second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset Metric ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ D.2 Supplementary Utility Results
The raw experiment results for MLE is shown in Table 10. The results are also visualized in Figure 12.
Table 10: Performance comparison between different models in terms of MLE. Classification tasks are evaluated by weighted AUC ROC scores, and regression tasks are evaluated by $R^{2}$ scores. The last row shows the average relative error versus real. Best scores and second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset ML real ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ D.3 Supplementary Privacy Results
Dataset ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ D.4 RDD Evaluation Results
Table 12: Performance comparison between different models in terms of RDD. AUC-ROC scores are reported, the smaller values mean that the synthetic records are harder to be detected. Best scores and second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset ML ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ D.5 Computation Time
Dataset Action ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ Appendix E Comparison to Baselines without Public Code
There are two recent tabular data generation works, TabMT (Gulati & Roysdon, 2024) and TabDiff (Shi et al., 2024), that have not published their code by the time of the composition of this paper. Unfortunately, although all papers report MLE scores, which can be a basis of comparison, the values from these papers cannot be directly taken for comparison because of the different settings of evaluation, including datasets, splits, ML model, and evaluation metric. However, fortunately, these papers report the performance of a few other common baselines and report the performance of the real data too. Thus, we use the reported values in these papers as indirect evidence for comparing these models’ performances. The values are shown in Table 15. While we acknowledge that experiments are the most solid way to showcase superior performance, this section mainly serves as a temporary qualitative analysis when these works’ code is not public yet.

### v5

@@ 4.5 Efficiency
Meanwhile, compared to baseline auto-regressive transformers, TabTreeFormer demonstrates a significant advantage in terms of training time on large datasets (see Table 5) and generation time on datasets with a large number of features (see Fig. 9), which is consistent with our expectation due to the smaller model size and sampling without rejection. Raw training and generation time records can be found in Table 13 in Appendix D.5.
@@ 4.6 Ablation Study
Table 6: Design component ablation. Relative score change when the corresponding component in TTF-S is taken out.
@@
We study the necessity of each building block in TabTreeFormer. It can be seen from Table 6 that TabTreeFormer experiences different degrees of noticeable performance degradation when we eliminate the tree-based model (TBM), dual-quantization tokenization (DQT), and OCEL, respectively, which verifies the indispensability of all building blocks. More results are presented in Appendix D.6.
@@ 5 Conclusion
In this paper, we propose TabTreeFormer, a hybrid tree-transformer for high-quality tabular data generation. TabTreeFormer inherits tabular-specific inductive biases from a tree-based model and captures multimodal distribution for optimized numeric data representations via a dual-quantization tokenizer. We comprehensively evaluate the performance of TabTreeFormer and show that it demonstrates outstanding fidelity, utility, privacy, and efficiency with significantly smaller model sizes and computation time.
Acknolwedgements
We would like to acknowledge Rishav Chourasia for helpful discussions.
@@ A.2 Proof of Theorem 1
Express $\mathcal{L}_{\text{oce}}$ and $\mathcal{L}_{\text{ce}}$ using $\mathbf{p}$, we have
$\mathcal{L}_{\text{ce}}(\mathbf{z},t)$ $=-\log p_{t}$ (6)
$\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ $=-\log\frac{w_{tt}e^{z_{t}}}{\sum_{i=1}^{V}w_{ti}e^{z_{i}}}$
$=-\log\frac{w_{tt}\cdot\frac{e^{z_{t}}}{\sum_{i=1}^{V}e^{z_{i}}}}{\sum_{i=1}^{V}w_{ti}\cdot\frac{e^{z_{i}}}{\sum_{j=1}^{V}e^{z_{j}}}}$
$=-\log\frac{w_{tt}p_{t}}{\sum_{i=1}^{V}w_{ti}p_{i}}$
$=-\log\frac{w_{tt}p_{t}}{D(\mathbf{z})}$ (7)
By $\mathcal{L}_{\text{ce}}(\mathbf{z},t)=\mathcal{L}_{\text{ce}}(\mathbf{z}^{*},t)$, we have $p_{t}=p_{t}^{*}$. Given $D(\mathbf{z})<D(\mathbf{z}^{*})$, we must have $\mathcal{L}_{\text{oce}}(\mathbf{z},t)<\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$.
@@ C.3 Baseline Implementation
We use the Synthcity (Qian et al., 2023) implementation for ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), GReaT (Borisov et al., 2023), TabDDPM (Kotelnikov et al., 2023), and TVAE (Xu et al., 2019). We also include CTAB-GAN+ (Zhao et al., 2024), REaLTabFormer (Solatorio & Dupriez, 2023), and TabSyn (Zhang et al., 2024) using their official code on GitHub or public SDK.
@@ D.1 Supplementary Fidelity Results
Table 9: Performance comparison between different models in terms of fidelity. The closer the values are to 1, the better the fidelity is. The best scores and the second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset Metric ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ D.2 Supplementary Utility Results
The raw experiment results for MLE is shown in Table 10. The results are also visualized in Figure 12.Out of the 30 MLE scores, TabTreeFormer-M or -L outperforms all other baselines in 20 cases (2/3), and they both outperform all other baselines in 16 cases ($\geq$1/2).
Table 10: Performance comparison between different models in terms of MLE. Classification tasks are evaluated by weighted AUC ROC scores, and regression tasks are evaluated by $R^{2}$ scores. The last row shows the average relative error versus real. The best scores and the second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset ML real ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ D.3 Supplementary Privacy Results
Dataset ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ D.4 RDD Evaluation Results
Table 12: Performance comparison between different models in terms of RDD. AUC-ROC scores are reported, the smaller values mean that the synthetic records are harder to be detected. The best scores and the second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset ML ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ D.5 Computation Time
Dataset Action ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
@@ Appendix E Comparison to Baselines without Public Code
There are two recent tabular data generation works, TabMT (Gulati & Roysdon, 2023) and TabDiff (Shi et al., 2024), that have not published their code by the time of the composition of this paper. Unfortunately, although all papers report MLE scores, which can be a basis of comparison, the values from these papers cannot be directly taken for comparison because of the different settings of evaluation, including datasets, splits, ML model, and evaluation metric. However, fortunately, these papers report the performance of a few other common baselines and report the performance of the real data too. Thus, we use the reported values in these papers as indirect evidence for comparing these models’ performances. The values are shown in Table 15. While we acknowledge that experiments are the most solid way to showcase superior performance, this section mainly serves as a temporary qualitative analysis when these works’ code is not public yet.